what percentage of workers walk to work?
Drive alone 7,862,418
carpool 1,120,982
take public transit 144,959
walk 138,496
other 149,641
work at home 346,626

Answer by jim_thompson5910(35256)

First add up all of the workers:
7,862,418+1,120,982+144,959+138,496+149,641+346,626=9,763,122

So there are a total of 9,763,122 workers in this sample


Now to find the percentage of workers who walk to work, simply divide the number of workers who walk by the total number of workers

138,496/9,763,122 = 0.014186


Now multiply by 100 to get a percentage

0.014186*100= 1.4186%


So about 1.42% of the workers walk to work